The graph of the parametric equations is called a cycloid.

x=θ-sinθ and y=1-cosθ for 0≤θ≤2π

(a) find dy/dx

(b) find an equation of the tangent to the cycloid at the point where θ=π/3

(c) at what point is the tangent horizontal?

(d) graph the cycloid and the tangent lines in parts (b) and (c)



1
Expert's answer
2021-11-18T09:17:26-0500

(a) "\frac{dy}{dx}=\frac{\frac{dy}{d\theta}}{\frac{dx}{d\theta}}"


"y=1-cos\theta ,\ \frac{dy}{d\theta}=sin\theta"


"x=\theta-sin\theta,\" "\frac{dx}{d\theta}=1-cos\theta"


"therefore \ \frac{dy}{dx}=\frac{sin\theta}{1-cos\theta}"



(b)Equation of tangent line to the cycloid

"when \ \theta=\frac{\pi}{3}\ we \ have"


"x=r(\frac{\pi}{3}-sin\frac{\pi}{3})=r(\frac{\pi}{3}-\frac{\sqrt{3}}{2})"


"y=r(1-cos\frac{\pi}{3})=\frac{r}{2}"

and"\frac{dy}{dx}=\frac{sin(\frac{\pi}{3})}{1-cos\frac{\pi}{3}}=\frac{\frac{\sqrt{3}}{2}}{1-\frac{1}{2}}=\sqrt{3}"

The slope of the tangent is "\sqrt{3}"


its equation is "y-\frac{r}{2}=\sqrt{3}(x-\frac{r\pi}{3}+\frac{r\sqrt{3}}{3})\ or \sqrt{3}x-y=r(\frac{\pi}{\sqrt{3}}-2)"



(c)The tangent is horizontal when "\frac{dy}{dx}=0"


This occurs when sin θ = 0 and 1 – cos θ ≠ 0,

that is, θ = (2n – 1)π, n an integer.


The corresponding point on the cycloid is

((2n – 1)π, 2).

When θ = 2nπ, both dx/dθ and dy/dθ are 0.



(d)Graph